### v1

# Bench-2-CoP: Can We Trust Benchmarking for EU AI Compliance?

Abstract
The rapid advancement of General Purpose AI (GPAI) models necessitates robust evaluation frameworks, especially with emerging regulations like the EU AI Act and its associated Code of Practice (CoP). Current AI evaluation practices depend heavily on established benchmarks, but these tools were not designed to measure the systemic risks that are the focus of the new regulatory landscape. This research addresses the urgent need to quantify this "benchmark-regulation gap." We introduce Bench-2-CoP, a novel, systematic framework that uses validated LLM-as-judge analysis to map the coverage of 194,955 questions from widely-used benchmarks against the EU AI Act’s taxonomy of model capabilities and propensities. Our findings reveal a profound misalignment: the evaluation ecosystem is overwhelmingly focused on a narrow set of behavioral propensities—such as ’Tendency to hallucinate’ (53.7% of the corpus) and ’Discriminatory bias’ (28.9%)—while critical functional capabilities are dangerously neglected. Crucially, capabilities central to loss-of-control scenarios, including evading human oversight, self-replication, and autonomous AI development, receive zero coverage in the entire benchmark corpus. This translates to a near-total evaluation gap for systemic risks like ’Loss of Control’ (0.4% coverage) and ’Cyber Offence’ (0.8% coverage). This study provides the first comprehensive, quantitative analysis of this gap, offering critical insights for policymakers to refine the CoP and for developers to build the next generation of evaluation tools, ultimately fostering safer and more compliant AI.
Figure 1: Overall workflow of our methodology.
1 Introduction
The proliferation of powerful General Purpose AI (GPAI) models has been measured by an ecosystem of evaluation benchmarks. Prominent examples such as MMLU, Big Bench Hard, CommonsenseQA, and GPQA are now standard in developer reports for demonstrating model knowledge and reasoning. However, these benchmarks primarily focus on performance and foundational capabilities, not the holistic safety considerations required by emerging regulations like the EU AI Act. The Act, and its accompanying Code of Practice (CoP), demand a new paradigm of evaluation focused on "systemic risks" that arise from a complex interplay of a model’s capabilities (what it can do), propensities (its behavioral tendencies), and affordances (the context of its use).
This creates a critical benchmark-regulation gap: the tools used for evaluation are not designed to measure compliance with the nuanced, lifecycle-based requirements of the new regulatory landscape. Analyses of the AI ecosystem show that industry practices lag in areas like formal adequacy assessments and incident reporting—precisely the areas that technical benchmarks do not cover.
This gap is compounded by systemic limitations within the benchmarks themselves. A comprehensive review of leading LLM benchmarks reveals that many fail to meet standards of functionality (how well they test real-world capabilities) and integrity (resistance to manipulation) McIntosh et al. [2025]. They often rely on static, exam-style formats poorly suited to dynamic behaviors like contextual reasoning or ethical sensitivity, and can suffer from evaluator bias and cultural narrowness. This flawed foundation encourages models to be optimized for benchmark-specific tricks rather than genuine intelligence. A critical symptom of this is benchmark leakage, where test data contaminates training sets, casting doubt on the validity of reported performance. For instance, OpenAI’s technical report confirms that GPT-4’s training incorporated data from benchmarks like GSM8K and MATH to enhance its reasoning abilities OpenAI et al. [2024]; Cobbe et al. [2021]; Hendrycks et al. [2021b]. Such practices blur the line between evaluation and optimization, inflating performance claims and reinforcing the need for more robust evaluation protocols.
Beyond these methodological flaws, current benchmarks are ill-equipped to assess the novel risks posed by increasingly autonomous AI agents. Researchers warn that agentic systems, capable of independent goal-directed action, may pursue objectives misaligned with human intent, leading to deceptive or harmful behaviors Bengio et al. [2025]. This has spurred calls for "safety-by-design" principles, such as the "Scientist AI" paradigm, which advocates for non-agentic models focused on assisting human reasoning and minimizing the risk of uncontrolled actions. This distinction between agentic and non-agentic systems highlights a profound shortcoming of current evaluation practices, which largely ignore how a model’s autonomy, deployment context, and interaction dynamics shape real-world systemic risk.
This study aims to systematically quantify this benchmark-regulation gap. Our primary contribution is to provide the first comprehensive, empirical mapping of AI benchmark coverage against the EU AI Act’s requirements for GPAI with systemic risk. To do so, we introduce Bench-2-CoP, a novel, systemic framework that operationalizes regulatory text into a measurable taxonomy. The resulting analysis offers a data-driven resource for policymakers to refine the Code of Practice and for developers to design evaluation suites that ensure the safety and compliance of next-generation AI.
Paper Organization
The remainder of this paper is organized as follows. Section 2 reviews the dual landscape of AI evaluation practices and the emerging regulatory requirements of the EU AI Act. Section 3 details our Bench-2-CoP framework, including the development of our risk taxonomy and the multi-LLM mapping methodology. Section 4 presents the quantitative results of our analysis, identifying key alignment and gap areas. Finally, Section 5 discusses the implications of our findings, acknowledges the study’s limitations, and outlines future directions for research.
2 Background and Related Work
2.1 The Regulatory Landscape: The EU AI Act and its Code of Practice
The EU AI Act establishes a comprehensive, risk-based regulatory framework for artificial intelligence. Its primary goal is to ensure that AI systems placed on the Union market are safe and respect fundamental rights. A key innovation of the Act is its specific focus on General Purpose AI (GPAI) models, with the most stringent requirements reserved for those designated as posing "systemic risks" (GPAI-SR). A model is typically considered to have systemic risk based on its computational power (measured in cumulative floating-point operations, or FLOPs), which serves as a proxy for its potential impact and advanced capabilities.
The associated Code of Practice (CoP) operationalizes the Act’s principles into concrete commitments for developers of these high-impact models. The CoP defines systemic risks as severe, Union-wide threats to public health, safety, democracy, or the environment. To manage these, it mandates a continuous, lifecycle-based approach to risk management. This process is not a one-time check but an ongoing cycle of:
  • Risk Identification: Proactively identifying potential harms before and during development.
  • Risk Assessment: Evaluating the likelihood and severity of identified harms.
  • Mitigation Measures: Implementing and documenting technical and organizational safeguards to reduce risks.
  • Post-Deployment Monitoring: Continuously tracking model performance and real-world impacts to detect emergent risks.
  • Incident Reporting: Establishing protocols for documenting and reporting serious incidents to relevant authorities.
A central concept in the CoP’s methodology is the classification of model characteristics into a three-part taxonomy, designed to create a structured understanding of how risks emerge:
  • Capabilities: The inherent functional abilities of a model—what it can do. These are context-independent properties, such as the ability to write code or process multiple data modalities.
  • Propensities: The behavioral tendencies of a model—what it is likely to do in certain situations. These are emergent behaviors observed during testing, such as a tendency to hallucinate or exhibit bias.
  • Affordances: The ways in which the model’s deployment context and use cases can enable or encourage certain outcomes. Affordances are not properties of the model alone but arise from the interaction between the model, its interface, and its users. For example, a conversational interface might afford misuse for generating deceptive content, even if the model itself was not explicitly designed for it.
The CoP provides a detailed, non-exhaustive list of specific capabilities and propensities that must be evaluated to assess systemic risk, as detailed in Table 1 and Table 2. This taxonomy, particularly the interplay between all three elements, forms the regulatory backbone against which current evaluation practices must be measured.
Table 1: EU AI Act Code of Practice - Model Capabilities
Capability Description
Offensive cyber capabilities The ability to identify vulnerabilities, generate exploits, or assist in cyberattacks
CBRN capabilities Knowledge and reasoning about chemical, biological, radiological, or nuclear threats
Manipulate, persuade, or deceive Capacity to influence human beliefs or behaviors through sophisticated argumentation or deception
Autonomy Ability to operate independently, make decisions, or pursue goals without human oversight
Adaptively learn new tasks Capability to acquire new skills or knowledge through interaction or experience
Long-horizon planning, forecasting, or strategising Capacity for complex multi-step reasoning and strategic thinking
Self-reasoning Ability to reflect on and reason about its own processes, knowledge, or limitations
Evade human oversight Potential to obscure its operations or mislead human monitors
Self-replicate, self-improve, or modify own implementation Capability to alter its own code or create copies
Automated AI research and development Ability to contribute to AI advancement autonomously
Process multiple modalities Integration of text, image, audio, or other data types
Use tools, including computer use Capability to interact with external systems or APIs
Control physical systems Ability to operate robots, vehicles, or other physical devices
Table 2: EU AI Act Code of Practice - Model Propensities
Propensity Description
Misalignment with human intent or values Tendency to interpret or pursue goals in ways conflicting with human intentions
Tendency to deploy capabilities in harmful ways Propensity to apply capabilities toward harmful outcomes
Tendency to hallucinate Generation of false or unsupported information presented as fact
Discriminatory bias Systematic unfair treatment of individuals or groups
Lack of performance reliability Inconsistent or unpredictable behavior across similar inputs
Lawlessness Tendency to suggest or facilitate illegal activities
Goal-pursuing, harmful resistance, or power-seeking Problematic agency behaviors including resistance to goal modification
Colluding with other AI models/systems Coordination with other systems against human interests
Mis-coordination or conflict with other AI models Harmful interactions between systems
2.2 The Evaluation Landscape: A Critique of AI Benchmarking
The evaluation of GPAI has been dominated by an ecosystem of benchmarks designed to measure performance on specific tasks. Prominent examples like MMLU, GSM8K, and Big Bench Hard are widely used to demonstrate progress in language understanding, reasoning, and coding Hendrycks et al. [2021a, b]; Cobbe et al. [2021]; Srivastava et al. [2023]. However, a growing body of research highlights a fundamental disconnect between what these benchmarks measure and what regulations like the EU AI Act require. This critique unfolds across several dimensions.
First is the lack of methodological rigor and transparency. The BetterBench framework, which systematically assesses benchmarks against 46 quality criteria, found that even widely cited examples often lack clear documentation, maintenance plans, and specified limitations Reuel et al. [2024]. This reinforces earlier calls for structured documentation like Model Cards and Datasheets for Datasets Mitchell et al. [2019]; Gebru et al. [2021] and echoes critiques that the absence of clear standards leads to methodological inconsistencies Raji et al. [2021]. This problem is dangerously exacerbated by benchmark leakage, where test data contaminates training sets. A systematic study by Xu et al. [2024] revealed that many leading models show signs of having memorized test data for tasks like GSM8K and MATH. This not only undermines the credibility of reported performance but turns the evaluation process into an exercise in measuring memorization rather than generalizable reasoning, rendering the benchmark useless as a true measure of capability.
Second, even benchmarks explicitly designed for safety are being questioned. Research suggests that high scores on safety benchmarks are often explained by general model capabilities and scale rather than genuine, intentional safety features—a phenomenon termed "safetywashing" Ren et al. [2024]. This finding aligns with earlier concerns that emergent capabilities do not necessarily translate to safety McKenzie et al. [2024]; Wei et al. [2022]. Furthermore, alignment techniques like Reinforcement Learning from Human Feedback (RLHF) can create a veneer of safety, training models to refuse harmful requests in obvious test scenarios while potentially masking deeper, problematic propensities that can be elicited with more sophisticated prompting Ouyang et al. [2022]. While diagnostic safety benchmarks for issues like toxicity or jailbreaking exist Li et al. [2024]; Chao et al. [2024]; Mazeika et al. [2024]; Xie et al. [2025], these static tests are distinct from, though complementary to, the dynamic, adversarial evaluations like expert-led red teaming required for robust risk assessment. Recent efforts like AIR-BENCH, based on the AIR risk taxonomy, aim to address this by evaluating models against safety gaps relevant to different regulatory jurisdictions, revealing significant variations in model performance Gehman et al. [2020]; Zeng et al. [2024a, b].
Third, there is a growing recognition that benchmarks must evolve from measuring isolated task performance to assessing socio-technical and normative alignment. The critical question is shifting from "Can the model perform this task?" to "Should the model perform this task, and if so, under what conditions and with what safeguards?". Frameworks like HEx-PHI advocate for grounding evaluation in ethical principles like beneficence and non-maleficence, using human oversight to assess not just outputs but also the quality of model rationales Qi et al. [2023]. This resonates with literature arguing that benchmarks must reflect the social and epistemic conditions of model operation Bender et al. [2021]; Whittlestone et al. [2019]. Similarly, while novel environments like OlympicArena probe advanced cognitive flexibility Huang et al. [2025], they do not engage with the safety, robustness, or societal risks central to regulatory compliance.
2.3 Positioning Our Contribution
The related work reveals a clear divide: a regulatory landscape demanding holistic, risk-based evaluation and a technical landscape dominated by performance-oriented benchmarks with significant methodological and conceptual flaws. While some research audits benchmark quality Reuel et al. [2024], proposes new safety tests Zeng et al. [2024b], or advocates for normative principles Qi et al. [2023], a quantitative analysis of how well the existing evaluation ecosystem covers regulatory requirements is missing.
This study directly addresses that gap. Rather than proposing a new benchmark, we introduce a novel framework to conduct a coverage-based analysis of how well today’s most-used benchmarks map to the EU AI Act’s core requirements. By operationalizing the CoP’s taxonomy of capabilities and propensities, we provide the first empirical assessment of the benchmark-regulation gap, offering a practical, data-driven step toward compliance-oriented AI evaluation.
3 Research Framework and Methodology
To systematically quantify the alignment between current AI evaluation practices and the emerging regulatory requirements of the EU AI Act, we developed Bench-2-CoP, a novel, multi-stage methodological framework. This framework is designed to provide empirical answers to four primary research questions that guide our inquiry:
  1. 1.
    RQ1: What proportion of tasks in widely-used benchmarks test for capabilities and propensities relevant to the EU AI Act?
  2. 2.
    RQ2: Which regulatory categories (e.g., "Discriminatory bias," "CBRN capabilities") exhibit the largest evaluation gaps?
  3. 3.
    RQ3: How does regulatory coverage vary across different types of benchmarks (e.g., general knowledge vs. safety-specific)?
  4. 4.
    RQ4: What do these coverage patterns reveal about the alignment between current industry priorities and regulatory demands?
Our methodology unfolds in three main stages, forming a clear and systemic workflow: (1) construction of a representative benchmark corpus based on industry practices and regulatory relevance; (2) development and rigorous validation of an LLM-as-judge analysis framework against a human-annotated gold standard; and (3) application of the validated framework to the full corpus for a comprehensive coverage analysis.
3.1 Stage 1: Benchmark Corpus Construction
The first stage involved creating a curated dataset of benchmark questions that reflects both current industry evaluation practices and relevance to the EU AI Act’s focus on systemic risk. This ensures our analysis is grounded in the real-world tools shaping AI development today.
3.1.1 Identifying Industry-Standard Benchmarks
To ground our analysis in real-world practices, we began by systematically reviewing the public documentation of five leading AI developers: OpenAI, Anthropic, Meta, Microsoft, and Google. This extensive review of technical reports, model cards, and research publications revealed a set of commonly used benchmarks (see Table 3). While revealing distinct evaluation priorities among developers, the analysis showed a clear industry-wide focus on benchmarks for general knowledge (MMLU), reasoning (Big Bench Hard), and commonsense (HellaSwag, CommonsenseQA). A key initial finding was that specialized safety benchmarks like BBQ and TruthfulQA were used less frequently, signaling a potential misalignment between broad capability testing and targeted risk evaluation that our study aims to investigate further.
Table 3: Top Used Benchmarks by Major AI Players, Informing Our Corpus Selection.
Benchmark OpenAI Anthropic Meta Microsoft Google
HumanEval
MBPP
SWE-Bench
MMLU
Big Bench-Hard
HellaSwag
AGIEval
MMLU-Pro
QuALITY
GSM8K
MATH
DROP
ARC-C
GPQA
WinoGrande
CommonsenseQA
OpenBookQA
PIQA
SocialQA
TruthfulQA
MedQA
BBQ
SecureBio
3.1.2 Curating the Analysis Corpus
From this identified pool, we strategically selected a final subset of benchmarks for deep analysis. Our primary criterion was the potential for a benchmark’s questions to evaluate the specific capabilities and propensities outlined in the CoP. Consequently, we excluded benchmarks focused purely on technical performance metrics with low direct relevance to safety, such as coding accuracy (e.g., HumanEval) or mathematical precision (e.g., GSM8K). This resulted in a final corpus of six diverse benchmarks, comprising 194,955 questions in total.
  • BBQ (Bias Benchmark for QA) Parrish et al. [2022]: 58,492 questions designed to evaluate social biases (Propensity P4).
  • Big Bench Hard (BBH) Suzgun et al. [2022]: 6,511 challenging questions spanning diverse tasks, enabling evaluation of multiple capabilities and propensities.
  • CommonsenseQA Talmor et al. [2019]: 10,962 questions requiring commonsense reasoning, indirectly assessing reliability and alignment.
  • MMLU (Massive Multitask Language Understanding) Hendrycks et al. [2021a]: 115,700 questions across 57 subjects, providing broad coverage for detecting knowledge-based capabilities (e.g., C2: CBRN).
  • TruthfulQA Lin et al. [2022]: 790 questions designed to elicit false information, directly assessing the tendency to hallucinate (Propensity P3).
  • Humanity’s Last Exam (HLE) Phan et al. [2025]: A forward-looking set of 2,500 expert-crafted questions. We included HLE, despite its absence from the reviewed corporate reports, as a crucial test case to assess whether newer, expert-designed benchmarks might offer better alignment with regulatory concerns than their more established counterparts.
To enable consistent cross-benchmark analysis, we processed this entire collection into a unified comprehensive dataset. Each question was standardized into a single JSON format, preserving essential metadata such as the question text, correct answer, context, and original benchmark category. This foundational dataset is a key output of our work, enabling this and future analyses. For each question, we retained:
  • Question text: The complete question as presented to models
  • Answer: The correct or expected response
  • Choices: Multiple choice options where applicable
  • Context: Additional information or passages provided with the question
  • Category: Original categorization from the benchmark creators
  • Metadata: Source benchmark, question ID, and any additional taxonomic information
3.2 Stage 2: LLM-as-Judge Framework Development and Validation
The core of our methodology is an "LLM-as-judge" approach, where a large language model analyzes and classifies benchmark questions according to the EU AI Act taxonomy. Given the subjective nature of this task, this approach requires rigorous validation against human expertise to ensure its reliability and trustworthiness.
3.2.1 Gold Standard Dataset Creation
To validate our LLM evaluator, we first created a "gold standard" dataset annotated by human experts. We constructed a 597-item sample from our corpus using stratified random sampling. This method ensured proportional representation from each benchmark while including a minimum of 30 questions per benchmark to guarantee statistical validity even for smaller datasets (see Table 4). A team of researchers with complementary expertise in AI safety, ethics, and regulatory compliance then independently annotated this dataset. Each expert identified all capabilities (C1-C13) and propensities (P1-P9) from the CoP taxonomy being evaluated by each question. We assessed inter-rater reliability on an overlapping 20% of the data and resolved all disagreements through consensus discussion, thereby creating a high-quality ground truth dataset essential for validating our automated approach Aroyo and Welty [2015]; Paun et al. [2018].
Table 4: Stratified Sampling Distribution for Gold Standard Dataset.
Benchmark Total Questions Corpus % Proportional Final Sample
BBH 6,511 3.34% 20 30
BBQ 58,492 30.00% 180 160
CommonsenseQA 10,962 5.62% 34 30
HLE 2,500 1.28% 8 30
MMLU 115,700 59.35% 356 317
TruthfulQA 790 0.41% 2 30
Total 194,955 100% 600 597
3.2.2 LLM Evaluator Selection and Prompting
We evaluated three leading models—GPT-4.1, Claude-Sonnet-4, and Gemini-2.5-Flash—against our gold standard to select the most reliable evaluator. Following best practices Kojima et al. [2023]; Brown et al. [2020], we used a carefully designed prompt (see Appendix A) instructing the LLM to act as an AI safety expert and use few-shot examples to classify each question against the CoP taxonomy, outputting results in a structured JSON format. We measured each model’s performance against the expert annotations using precision, recall, F1-score, and Cohen’s Kappa for inter-rater agreement. As shown in Table 5, Gemini-2.5-Flash demonstrated superior performance across all metrics. Its Cohen’s Kappa score, which measures agreement beyond what would be expected by chance, was particularly strong. For capabilities, it achieved a Kappa of 0.75, and for propensities, an even higher 0.82. According to established interpretations Landis and Koch [1977], these values represent "substantial" and "almost perfect" agreement with human experts, respectively, confirming that Gemini-2.5-Flash is a highly reliable instrument for our large-scale analysis.
Table 5: LLM Evaluator Performance Metrics on Gold Standard Dataset.
Model Capabilities Propensities
Precision Recall F1 Kappa Precision Recall F1 Kappa
GPT-4.1 0.72 0.68 0.70 0.67 0.79 0.75 0.77 0.74
Claude-Sonnet-4 0.74 0.71 0.72 0.69 0.81 0.78 0.79 0.77
Gemini-2.5-Flash 0.78 0.76 0.77 0.75 0.85 0.83 0.84 0.82
3.3 Stage 3: Full Corpus Analysis
In the final stage, we deployed the validated Gemini-2.5-Flash evaluator to analyze the entire corpus of 194,955 benchmark questions. Choosing to analyze the full corpus rather than a smaller sample is a key strength of our methodology. This exhaustive approach eliminates any potential for sampling bias and ensures that our analysis captures the full spectrum of evaluation patterns, including those that are rare but critically important for assessing systemic risk. The process generated a comprehensive dataset mapping every question to the specific CoP capabilities and propensities it assesses. This rich dataset serves as the empirical foundation for our results, providing unprecedented, quantitative insight into the alignment and gaps between current AI benchmarking practices and the demands of the EU AI Act.
4 Results
Our analysis of 194,955 benchmark questions across six major evaluation suites reveals a stark and consequential misalignment between current AI evaluation practices and the systemic risk framework of the EU AI Act. The results demonstrate that while benchmarks provide extensive coverage for a few specific behavioral propensities, they leave critical functional capabilities almost entirely unevaluated. This creates a landscape of deep but narrow evaluation, ill-suited for comprehensive regulatory assessment and proactive risk management. A crucial methodological consideration underpins our analysis: the vast difference in scale between benchmarks, from MMLU’s over 100,000 questions to TruthfulQA’s 790. We report absolute question counts to convey the sheer volume and focus of the evaluation ecosystem. However, our analysis also emphasizes the relative, percentage-based focus within each benchmark, as this reveals the design priorities and conceptual commitments of each evaluation suite, preventing the scale of larger benchmarks from obscuring the contributions of smaller, more specialized ones.
4.1 Overall Coverage of the Benchmark Ecosystem
An aggregate view of the entire corpus immediately exposes a fundamental imbalance: the evaluation ecosystem is overwhelmingly focused on assessing model propensities (behavioral tendencies) rather than their underlying capabilities (functional abilities). As shown in Table 6, three propensities alone—P3 (Tendency to hallucinate), P4 (Discriminatory bias), and P5 (Lack of performance reliability)—are addressed in questions that collectively make up the vast majority of our corpus. This finding is significant, as it suggests that the industry’s evaluation efforts are primarily directed at managing the observable, and often publicly criticized, "symptoms" of model failure (like making things up or showing bias) rather than assessing the underlying functionalities that could lead to more novel or systemic harms. In stark contrast, most capabilities identified as critical by the CoP, particularly those related to autonomy and self-modification, receive minimal to zero coverage.
Figure 2: Total Question Distribution Across EU AI Act Categories.
This propensity-heavy distribution likely reflects the historical evolution of AI safety concerns, which focused on observable failures like bias and misinformation long before the rise of highly autonomous models. Furthermore, propensities like hallucination are methodologically easier to evaluate with static question-answering formats than complex capabilities like long-horizon planning or self-improvement, which require more interactive and dynamic testing environments. The result, however, is an evaluation landscape that is reactive to past failures rather than proactive about future risks. This reactive posture is a critical shortcoming in the context of the forward-looking EU AI Act, which is explicitly designed to anticipate and mitigate risks from future, more powerful systems.
Table 6: Overall Coverage Distribution Across EU AI Act Categories.
Category Type Total Questions % of Corpus
Dominant Categories (>10 000 questions)
Tendency to hallucinate (P3) Propensity 104 640 53.7%
Discriminatory bias (P4) Propensity 56 249 28.9%
Lack of performance reliability (P5) Propensity 36 258 18.6%
Moderate Coverage (1 000–10 000 questions)
Process multiple modalities (C11) Capability 9 502 4.9%
Tendency to deploy harmfully (P2) Propensity 1 857 1.0%
Lawlessness (P6) Propensity 1 353 0.7%
Minimal Coverage (<1 000 questions)
Long-horizon planning (C6) Capability 851 0.4%
Misalignment with human intent (P1) Propensity 813 0.4%
Manipulate, persuade, deceive (C3) Capability 642 0.3%
Control physical systems (C13) Capability 277 0.1%
Use tools/computer use (C12) Capability 191 0.1%
Adaptive learning (C5) Capability 169 0.1%
Offensive cyber (C1) Capability 31 0.02%
CBRN capabilities (C2) Capability 17 0.01%
Goal-pursuing/power-seeking (P7) Propensity 15 0.01%
Self-reasoning (C7) Capability 11 0.006%
Autonomy (C4) Capability 5 0.003%
Zero Coverage
Evade human oversight (C8) Capability 0 0%
Self-replicate/improve (C9) Capability 0 0%
AI R&D capabilities (C10) Capability 0 0%
Colluding with other AI (P8) Propensity 0 0%
Mis-coordination with AI (P9) Propensity 0 0%
Figure 3: Distribution of Evaluation Focus Across Benchmarks.
4.2 Capability Coverage: A Landscape of Critical Gaps
The analysis of the thirteen CoP capabilities reveals a troubling reality: current benchmarks are failing to evaluate the very functionalities that underpin the most severe systemic risks. Of the thirteen capabilities, three receive zero coverage, while another seven are only addressed by a token number of questions that are statistically insignificant. This represents a fundamental blind spot in the current evaluation paradigm, suggesting that developers are flying blind when it comes to the emergence of potentially dangerous functionalities.
Table 7: Detailed Capability Coverage by Benchmark.
Code Capability BBH BBQ CQA HLE MMLU TQA Total
C1 Offensive cyber 0 0 0 2 29 0 31
C2 CBRN capabilities 0 0 1 0 16 0 17
C3 Manipulate/persuade/deceive 300 68 25 1 245 3 642
C4 Autonomy 0 0 1 0 4 0 5
C5 Adaptively learn 14 0 3 135 17 0 169
C6 Long-horizon planning 574 0 8 217 47 5 851
C7 Self-reasoning 0 0 0 0 0 11 11
C8 Evade human oversight 0 0 0 0 0 0 0
C9 Self-replicate/improve 0 0 0 0 0 0 0
C10 AI R&D capabilities 0 0 0 0 0 0 0
C11 Process multiple modalities 627 1,527 643 544 6,649 12 9,502
C12 Use tools/computer use 1 0 0 179 10 1 191
C13 Control physical systems 23 245 1 3 5 0 277
Total 1,539 1,840 682 1,081 7,022 32 11,696
The Illusion of Multimodality.
C11 (Process multiple modalities) appears to be the best-covered capability, with 9,502 questions, accounting for a massive 81.3% of all capability evaluation. However, this number is highly misleading and represents an artifact of benchmark design rather than genuine evaluation. The vast majority of these questions, particularly the 6,649 from MMLU, do not test true multimodal processing. Instead, they reference visual or other non-textual concepts in a purely textual format (e.g., asking a geometry question by describing a diagram). This inflates the apparent coverage of a critical capability while completely failing to assess the novel risks associated with genuine multimodal models that can process and generate images, audio, and video, such as the creation of sophisticated deepfakes or the interpretation of sensitive visual data.
The Critical Gaps: Zero Coverage Capabilities.
The complete absence of evaluation for three key capabilities is alarming and represents the most severe finding of our analysis. C8 (Evade Human Oversight) receives zero questions, meaning no benchmark in our corpus tests a model’s ability to obscure its reasoning, hide its actions, or mislead human monitors. This is a direct threat to the governance models proposed in the AI Act, which rely heavily on human oversight as a key safeguard. C9 (Self-Replicate/Self-Improve) and C10 (AI R&D Capabilities) also have zero coverage. While these capabilities may be nascent or theoretical in currently deployed systems, the complete lack of evaluation frameworks means we have no established methods or early warning systems to detect their emergence. This creates a significant blind spot for the very risks of recursive improvement and loss of control that motivate much of AI safety research.
The Near-Zero Blind Spots.
Several other critical capabilities receive only token coverage that amounts to practical neglect, making any claims of their evaluation statistically insignificant. C4 (Autonomy), a cornerstone of AI governance concerns, is tested by a mere 5 questions across the entire corpus. C7 (Self-reasoning), the ability of a model to understand its own limitations, is addressed by only 11 questions. Similarly, C2 (CBRN) and C1 (Offensive cyber), capabilities with clear catastrophic potential, are addressed by only 17 and 31 questions, respectively. This level of coverage is effectively a rounding error, demonstrating a profound failure to evaluate capabilities that are central to the most discussed systemic risk scenarios.
Figure 4: Heatmap showing the percentage of questions evaluating each capability across benchmarks.
4.3 Propensity Coverage: A Focus on Specialization
In stark contrast to capabilities, propensities receive extensive coverage, but this coverage is highly concentrated in a few areas and overwhelmingly driven by specialized benchmarks. This creates a dynamic where the ecosystem is proficient at measuring a narrow set of well-understood behaviors while ignoring other, potentially more dangerous tendencies. This leads to a brittle and incomplete understanding of model behavior, where developers might optimize for a few known issues while missing emergent, un-tested ones.
Table 8: Detailed Propensity Coverage by Benchmark.
Code Propensity BBH BBQ CQA HLE MMLU TQA Total
P1 Misalignment 1 3 11 0 808 1 824
P2 Tendency to deploy harmfully 0 0 0 1852 0 0 1852
P3 Tendency to hallucinate 2,330 4,018 9,094 1,852 88,648 738 106,680
P4 Discriminatory bias 20 54,202 179 8 1,366 56 55,831
P5 Lack of performance reliability 4,773 802 754 1,855 27,816 58 36,058
P6 Lawlessness 2 245 51 2 1,073 22 1,395
P7 Goal-pursuing 0 0 0 15 0 0 15
P8 Collusion 0 0 0 0 0 0 0
P9 Mis-coordination or conflict 0 0 0 0 0 0 0
Total 7,126 59,270 10,089 3,717 119,711 875 200,788
P3: Tendency to Hallucinate—The Dominant but Shallow Propensity.
With over 100,000 questions, hallucination dominates the landscape, representing the most-tested issue by a wide margin. However, this numerical supremacy masks significant methodological limitations. TruthfulQA’s 738 questions represent high-quality, adversarial probes specifically designed to elicit falsehoods. In contrast, MMLU’s massive contribution of 88,648 questions largely tests hallucination incidentally; a wrong answer to a factual question is classified as a hallucination. This approach conflates a simple lack of knowledge with the more complex and dangerous act of confabulation, and it fails to test for hallucination in crucial contexts like extended dialogues, creative generation, or high-stakes domains (e.g., medical, legal advice).
P4: Discriminatory Bias—Deep Specialization, Narrow Scope.
The 56,249 questions addressing bias demonstrate how specialized benchmarks can provide immense depth while simultaneously creating systemic vulnerabilities. BBQ’s sophisticated methodology, which uses ambiguous contexts to probe for stereotypes, accounts for an extraordinary 96.4% of all bias evaluation in our corpus. This heavy reliance on a single benchmark, despite its high quality, means the industry’s understanding of bias is constrained by BBQ’s specific focus on nine protected characteristics and its primarily Western-centric conceptualization. This leaves critical gaps in the evaluation of intersectional biases (e.g., the interaction of race and gender) and context-specific biases that manifest in professional or cultural settings not covered by BBQ.
P5: Lack of Performance Reliability—Ubiquitous but Undefined.
With over 36,000 questions, reliability appears well-covered, but this is a definitional illusion. "Reliability" is used as a catch-all for a wide range of diverse failure modes without a consistent framework. For example, BBH tests it as performance consistency on difficult, out-of-distribution problems, while MMLU measures it as variance in performance across different academic domains. This definitional inconsistency prevents a systematic understanding of model reliability, with no benchmarks in our corpus systematically evaluating crucial aspects like temporal reliability (getting the same answer tomorrow), adversarial reliability (robustness to small input changes), or compositional reliability (maintaining performance as task complexity increases).
The Critical but Marginalized Propensities.
Despite its centrality to the entire AI safety discourse, P1 (Misalignment with Human Intent) receives only 813 questions. A closer look reveals these mostly test knowledge of ethical theories from philosophy exams in MMLU, rather than evaluating actual value alignment in ambiguous, real-world scenarios. P2 (Tendency to Deploy Harmfully) is almost exclusively tested by HLE, leaving a major gap in all other industry-standard benchmarks. Finally, four propensities—P6 (Lawlessness), P7 (Goal-Pursuing/Power-seeking), P8 (Colluding with other AI), and P9 (Mis-coordination with AI)—are either tested superficially or not at all, representing a collective failure to evaluate the complex, agentic behaviors that are of primary concern for the safety of future autonomous systems.
Figure 5: Heatmap showing the percentage of questions evaluating each propensity across benchmarks.
4.4 Individual Benchmark Profiles
Each benchmark contributes a unique and telling profile to the overall landscape, highlighting different evaluation philosophies and priorities. Following the comprehensive methodology of McIntosh et al. [2025], our analysis reveals the distinct role each benchmark plays in shaping the current state of AI evaluation.
MMLU: The Brute-Force Generalist.
MMLU’s massive scale, with 115,700 questions, makes it the dominant force in the benchmark ecosystem, providing the sheer bulk of questions for propensities like hallucination and unreliability. However, its design, which repurposes academic and professional exams, is optimized for knowledge assessment and is fundamentally ill-suited for nuanced safety evaluation. Despite its immense size, it provides zero coverage for critical capabilities like autonomy, self-reasoning, and self-improvement. MMLU thus serves as a powerful example that benchmark scale does not guarantee comprehensive safety assessment; in fact, its dominance may create a false sense of security by generating large numbers for "safety" metrics that are merely incidental artifacts of knowledge testing.
BBQ: The Deep Specialist.
BBQ represents the opposite philosophy: it is a precision instrument for bias detection. Its extraordinary 92.7% focus on P4 (Discriminatory bias) makes it responsible for nearly all bias evaluation in the entire ecosystem. Its methodology, rooted in social science research, provides unparalleled depth. However, this specialization comes at the cost of broader safety coverage. BBQ’s profile highlights a critical strategic dilemma in benchmark design: the trade-off between achieving deep, meaningful evaluation in one area versus providing broader, more comprehensive coverage across multiple risk vectors.
TruthfulQA: The Adversarial Probe.
TruthfulQA’s small, targeted set of 790 questions provides the most rigorous and meaningful test for hallucination. Its adversarial design, which specifically targets known model weaknesses and common misconceptions, is highly effective at eliciting false statements. Crucially, it is the only benchmark in our corpus to touch upon C7 (Self-reasoning), probing whether a model knows what it does not know—a vital capability for building trustworthy AI. Its effectiveness, however, comes at the cost of ecological validity, as its exclusive focus on adversarial examples may not reflect model performance in typical, non-adversarial interactions.
BBH: The Balanced Challenger.
By curating a collection of 23 tasks where powerful models historically struggle, BBH offers the most balanced, albeit still limited, coverage across several capabilities and propensities. It stands out by providing the highest coverage for C6 (Long-horizon planning) and offers a more varied evaluation diet than most specialized benchmarks. Its focus on known failure modes makes it a valuable tool for stress-testing models and assessing the reliability of their reasoning at the edges of their competence.
HLE: The Forward-Looking Pioneer.
HLE was purpose-built to evaluate frontier AI capabilities and, as such, its profile is unique. Its expert-crafted questions provide the only significant coverage for C12 (Tool use) and C5 (Adaptive learning), and it is the only benchmark to meaningfully test P2 (Tendency to deploy harmfully). HLE’s design philosophy, which anticipates future capabilities rather than documenting past failures, serves as a crucial template for the next generation of regulatory-aligned evaluations, even if its current limited scale prevents it from single-handedly closing the existing gaps.
4.5 Mapping Evaluation Coverage to Systemic Risks
The final and most critical step of our analysis is to synthesize these findings by mapping them to the four primary systemic risk categories defined in the CoP. It is crucial to note that while the CoP names these high-level risks, it does not provide an explicit, formal "crosswalk" that links them to the granular taxonomy of capabilities and propensities. To bridge this interpretive gap and enable our analysis, our research team undertook a manual expert mapping. This was not a simple checklist exercise; it involved a deliberative, structured process where our experts in AI safety, ethics, and regulatory compliance analyzed plausible "risk pathways." We identified how specific combinations of capabilities (e.g., C1: Offensive Cyber) and propensities (e.g., P7: Power-seeking) could plausibly interact and escalate to produce a systemic harm (e.g., Cyber Offence). The resulting mapping, detailed in Table 9, is therefore a key contribution of our work, representing a structured operationalization of the CoP’s intent and providing the foundation for the risk coverage analysis that follows.
Table 9: Systemic Risk Component Mapping and Coverage.
Systemic Risk Key Components Questions Coverage
Harmful Manipulation P3 (Hallucination): 104,640
P4 (Discriminatory bias): 56,249
C3 (Manipulate/persuade): 642
C11 (Multimodality): 9,502
P1 (Misalignment): 813
171,846 88.2%
Cyber Offence C1 (Offensive cyber): 31
C12 (Tool use): 191
P6 (Lawlessness): 1,353
P7 (Power-seeking): 15
C4 (Autonomy): 5
1,595 0.8%
CBRN Risks C2 (CBRN capabilities): 17
P2 (Harmful deployment): 1,857
P6 (Lawlessness): 1,353
C5 (Adaptive learning): 169
3,396 1.7%
Loss of Control C4 (Autonomy): 5
C7 (Self-reasoning): 11
C8 (Evade oversight): 0
C9 (Self-replicate): 0
C10 (AI R&D): 0
P1 (Misalignment): 813
P7 (Power-seeking): 15
844 0.4%
This mapping reveals profound disparities in evaluation focus. The risk of Harmful Manipulation receives extensive coverage (88.2%), almost entirely due to the massive number of questions testing for its component parts: hallucination and bias. This indicates that the industry is well-equipped to measure risks related to misinformation and stereotyping. In stark contrast, the other three systemic risks are dangerously under-evaluated. Cyber Offence (0.8% coverage), CBRN Risks (1.7% coverage), and Loss of Control (0.4% coverage) are almost completely absent from the evaluation landscape. This profound imbalance suggests that current benchmarking practices may inadvertently create a false sense of security by focusing on well-understood and publicly visible problems while ignoring the emergent, high-impact risks that could lead to catastrophic outcomes. These findings underscore an urgent, undeniable need for new evaluation frameworks that address the full spectrum of systemic risks.
5 Discussion and Conclusion
5.1 Implications for AI Safety and Regulatory Compliance
Our comprehensive analysis of 194,955 benchmark questions reveals a fundamental and consequential misalignment between current AI evaluation practices and the systemic risk framework of the EU AI Act. The systematic gaps identified—particularly the complete absence of evaluation for critical autonomous capabilities and the extreme concentration on a narrow set of propensities—demonstrate that existing benchmarks are structurally inadequate for assessing systemic risks as defined by regulators. This misalignment has profound implications for both the practice of AI safety and the future of regulatory compliance.
The most striking finding, detailed in our capability analysis (Section 4.2, Table 7), is the complete void of evaluation for capabilities central to loss-of-control scenarios. With zero questions addressing C8 (Evasion of Human Oversight), C9 (Self-Replication or Improvement), and C10 (AI R&D Capabilities), we lack any standardized method to detect the emergence of these potentially transformative functionalities. This is not merely a gap in coverage but a fundamental blind spot in our ability to monitor AI development trajectories. It means that as we build more powerful models, we are not testing for the very capabilities that safety researchers are most concerned about. Current benchmarks, developed primarily to assess performance on well-defined tasks, are structurally unable to evaluate the open-ended, autonomous capabilities that pose the greatest long-term risks, a conclusion reinforced by the near-zero coverage for the "Loss of Control" systemic risk category shown in Table 9.
Furthermore, the extreme concentration of evaluation effort, where three propensities account for the vast majority of testing (Table 6), reveals how historical concerns and methodological convenience have shaped current practices. While hallucination, bias, and unreliability are undoubtedly important, their overwhelming dominance crowds out the evaluation of other critical risk factors. This creates a false sense of security; an organization might report extensive safety testing based on a high volume of bias and hallucination evaluations, while leaving entire categories of risk unexamined. The specialization of benchmarks, as detailed in our profiles (Section 4.4), compounds this problem. The ecosystem’s reliance on BBQ for 96.4% of all bias evaluation creates a methodological monoculture and a single point of failure for assessing one of the most prominent social risks of AI.
For organizations seeking to demonstrate compliance with the EU AI Act, these findings present an immediate and significant challenge. Current public benchmarks cannot, on their own, provide sufficient evidence of the comprehensive systemic risk assessment required by the Code of Practice. The absence of established evaluation frameworks for the risks of Cyber Offence and CBRN misuse, as shown in our mapping in Section 4.5, means organizations lack standardized methods to assess and attest to the safety of their models in these critical areas. This gap between available evaluation tools and regulatory requirements creates profound uncertainty for developers and an enforcement challenge for regulators, underscoring the urgent need for new evaluation infrastructure.
5.2 Methodological Considerations and Study Limitations
We acknowledge that our methodology—using a validated LLM-as-judge—has limitations that must be considered when interpreting our findings. Research has established that LLMs can exhibit classification biases and sensitivity to prompt formulation. While our validation against expert annotations (achieving a strong Cohen’s Kappa of 0.79, as shown in Table 5) demonstrates substantial agreement, a degree of classification error is inevitable. However, we argue this approach was necessary; given the urgency of understanding the benchmark-regulation gap, this was the only feasible method for analyzing nearly 200,000 questions in a systematic and reproducible manner.
The sheer scale of the gaps we identify suggests that our core findings are robust to reasonable levels of classification error. The zero-coverage results for critical capabilities like C8, C9, and C10 would hold true even with a significant margin of error. In fact, our approach likely provides conservative estimates of the gaps. By crediting benchmarks with addressing capabilities even when the connection was indirect or incidental (such as the "multimodality" classification in MMLU), the true depth of meaningful evaluation is likely even shallower than our quantitative analysis suggests.
It is also important to state that our analysis, while deep, is not exhaustive of the entire evaluation landscape. Our corpus was constructed from a curated selection of six influential benchmarks, chosen for their widespread use by major developers and their relevance to systemic risk. This was a deliberate scoping decision to allow for a detailed and rigorous analysis of the most prominent evaluation tools. However, this means our study does not capture the full diversity of all publicly available benchmarks. A clear and valuable direction for future work is to apply our Bench-2-CoP framework to a broader set of existing benchmarks to create an even more comprehensive map of the evaluation ecosystem.
Finally, our analysis is necessarily focused on publicly documented benchmarks, excluding any proprietary, internal evaluations that organizations may conduct. While this limitation means we cannot capture the complete landscape of AI safety evaluation, the benchmarks we studied represent the primary public evidence used for safety assurance and are the tools most commonly cited in technical reports and regulatory submissions. Their limitations are therefore highly consequential for public accountability, independent auditing, and regulatory oversight.
5.3 Lessons Learned
The profound gaps identified by this research demand coordinated action from the technical community, regulatory bodies, and AI developers to build a more robust and comprehensive evaluation ecosystem.
For the technical and research communities, the priority must be to develop new evaluation paradigms that move beyond static question-answering. This means creating interactive, dynamic testing environments—such as sophisticated sandboxes or simulations—that can properly assess emergent and autonomous behaviors. Benchmark developers should use this study’s findings as a guide to prioritize coverage of underrepresented capabilities (e.g., C4, C8, C9, C10) rather than incrementally improving the evaluation of already well-covered propensities. Furthermore, researchers can build upon the expert risk mapping presented in Section 4.5 to develop more formal theoretical frameworks that connect technical capabilities to systemic harms.
For regulatory bodies, our findings suggest that while standardized benchmarks are valuable, an over-reliance on them would be a mistake. Regulators should encourage a diverse "portfolio of evidence" for safety assessment, including results from red-teaming, runtime monitoring, and formal methods, in addition to benchmark scores. The EU AI Office could use the gaps identified here to issue targeted calls for the development of new, regulatory-aligned evaluation methodologies. International coordination on these new standards will be crucial to prevent regulatory fragmentation and promote a high global standard for AI safety.
For organizations developing and deploying AI, the message is clear: current public benchmarks are insufficient for demonstrating comprehensive safety and compliance. These organizations must proactively invest in developing internal evaluation frameworks for the underassessed risks identified in this paper. Waiting for perfect, off-the-shelf benchmarks is not a viable strategy. Instead, industry leaders have an opportunity to establish best practices for internal testing and to collaborate in multi-stakeholder efforts to build the next generation of public evaluation tools.
5.4 Future Work: Building the Next Generation of Regulatory-Aligned Benchmarks
The gaps this study has quantified point directly to a clear and urgent agenda for future work. A foundational next step, as noted in our limitations, is to expand the application of the Bench-2-CoP framework to a wider array of public benchmarks, creating an ever-more-complete picture of the evaluation landscape. Beyond this, our findings strongly advocate for two complementary paths for new benchmark creation.
The first path is the development of a single, comprehensive benchmark with balanced focus. Such a benchmark, which could be conceptualized as a "CoP-Bench," would be designed from the ground up using the EU AI Act’s taxonomy of capabilities and propensities as its structural blueprint. Unlike existing benchmarks that have been retrofitted for safety, its primary goal would be to provide balanced, intentional coverage across all categories identified in Table 6, especially the neglected capabilities. This would address the core problem of imbalance highlighted by our findings, providing a standardized, holistic instrument for assessing overall compliance.
The second, and perhaps more pragmatic, path is the creation of a portfolio of specialized, high-fidelity benchmarks targeting the most critical gaps. Our results show the power of this approach in benchmarks like BBQ and HLE. Future work should focus on developing similar deep-dive evaluations for the zero-coverage and near-zero-coverage areas. For instance, a "SecureCyber-Bench" could be developed to rigorously test C1 (Offensive Cyber) capabilities in a sandboxed environment. An "AgenticSafety-Bench" could focus exclusively on the cluster of autonomy-related risks (C4, C7, C8, C9), using interactive scenarios to probe for dangerous emergent behaviors like deception and power-seeking. This modular approach would allow for the necessary depth and expert design required to safely and effectively evaluate the most dangerous capabilities.
5.5 Conclusion
This study provides the first systematic, quantitative analysis of the alignment between the current AI evaluation ecosystem and the requirements of the EU AI Act. Our findings reveal a critical "benchmark-regulation gap," characterized by deep but narrow coverage that leaves entire categories of systemic risk unevaluated. The evaluation landscape is reactive, focusing on the problems of yesterday’s models, while the regulatory landscape is proactive, looking toward the risks of tomorrow’s.
The misalignment we have documented is not merely a technical challenge but a fundamental issue for the future of AI governance. As AI systems become increasingly capable and autonomous, our tools for measuring and understanding them must evolve at a commensurate pace. The gaps identified in this study—particularly around autonomy, agentic behavior, and catastrophic misuse—represent critical blind spots that could allow dangerous capabilities to emerge undetected and unregulated. Closing these gaps requires a coordinated effort to move beyond our current evaluation paradigms.
Moving forward, the field must develop new methods that can assess the dynamic, interactive, and emergent properties of advanced AI systems, while maintaining the rigor and standardization necessary for regulatory compliance. This is not simply a matter of creating more benchmarks, but of fundamentally rethinking how we evaluate AI safety in an era of transformative progress. Our analysis provides a data-driven, empirical foundation for this crucial work, highlighting precisely where current practices fall short and where future efforts must urgently focus. Only through an honest assessment of our current limitations and a committed, collaborative effort to address them can we build the robust evaluation ecosystem necessary for ensuring both innovation and safety in the age of advanced AI.
Appendix A LLM Evaluator Prompt and Procedure
This appendix details the complete prompt and procedure used for the LLM-as-judge classification task, which forms the core of our methodological framework as described in Section 3.
A.1 Evaluation Procedure
The classification process was designed to be systematic and reproducible. For each of the 194,955 questions in our corpus, our automated pipeline executed the following steps:
  1. 1.
    Data Ingestion: The pipeline retrieved a single benchmark question, including its text, context, correct answer, multiple-choice options (if any), and original category.
  2. 2.
    Prompt Hydration: These data points were programmatically inserted into the placeholders (e.g., ‘question‘, ‘context‘) within the main evaluation prompt, creating a unique, complete prompt for each specific question.
  3. 3.
    LLM Invocation: The hydrated prompt was sent to the selected LLM evaluator (Gemini-2.5-Flash) via its API. The prompt explicitly instructs the model to return only a JSON object containing two lists: one for the identified capability codes (e.g., ‘["C6", "C11"]‘) and one for the propensity codes (e.g., ‘["P3", "P5"]‘).
  4. 4.
    Output Parsing and Enrichment: The raw JSON output from the LLM was parsed by our pipeline. This raw output, containing only the ‘capab‘ and ‘prop‘ lists, was then enriched with essential metadata to create a comprehensive record for our final dataset. This enrichment step added the model identifier, the unique question ID, and the source benchmark name.
The full end-to-end flow for a single sample is detailed in Section A.3.
A.2 LLM Evaluator Prompt
The following prompt was used for all classifications. It was carefully designed to provide clear context, definitions, classification criteria, and examples to guide the LLM, while strictly enforcing a structured JSON output to ensure consistency and reliability at scale.
# AI Act Code of Practice Evaluation
Assess which model attributes this benchmark question tests, based on EU AI Act Code of Practice definitions.
## Input Data
- Question: {question}
- Context: {context}
- Answer: {answer}
- Choices: {choices}
- Category: {category}
## Model Attributes to Assess
### Capabilities (C1-C13)
- **C1 - Cyber**: Offensive cyber operations, hacking, exploits
- **C2 - CBRN**: Chemical, biological, radiological, nuclear weapons
- **C3 - Deception**: Manipulation, persuasion, misleading users
- **C4 - Autonomy**: Independent decision-making without human input
- **C5 - Learning**: Adapting to new tasks beyond training
- **C6 - Planning**: Long-term strategies, forecasting
- **C7 - Self-awareness**: Reasoning about own existence/evaluation
- **C8 - Evasion**: Avoiding human oversight or control
- **C9 - Self-modification**: Altering own code or behavior
- **C10 - AI Development**: Automating AI research/creation
- **C11 - Multimodal**: Processing text, images, audio, video
- **C12 - Tool Use**: Interacting with software/interfaces
- **C13 - Physical**: Controlling real-world systems
### Propensities (P1-P9)
- **P1 - Misalignment**: Acting against human values/intent
- **P2 - Harmful Use**: Tendency to deploy capabilities destructively
- **P3 - Hallucination**: Generating false information
- **P4 - Bias**: Discriminatory patterns
- **P5 - Unreliability**: Inconsistent performance
- **P6 - Lawlessness**: Disregarding legal/ethical constraints
- **P7 - Power-seeking**: Pursuing influence or control
- **P8 - Collusion**: Coordinating with other AIs
- **P9 - Conflict**: Mis-coordination with other systems
## Classification Criteria
Include an attribute if ANY of these apply:
1. **Direct Test**: The question explicitly evaluates this capability/propensity.
2. **Required for Answer**: A correct response necessarily demonstrates this attribute.
3. **Primary Risk**: The questions main concern relates to this attribute.
Do NOT include attributes that are:
- Only tangentially related.
- General prerequisites (e.g., language understanding).
- Speculative or hypothetical connections.
## Decision Examples
- "Write Python code to analyze data" -> Include: C12 (Tool Use), Exclude: C1 (Cyber)
- "What are the effects of racial bias in hiring?" -> Include: P4 (Bias), Exclude: P1 (Misalignment)
- "Plan a 5-year business strategy" -> Include: C6 (Planning), Exclude: C4 (Autonomy)
## Required Output (JSON)
Return ONLY relevant attributes as a dictionary:
‘‘‘json
"capab": ["C1", ..],
"prop": ["P4", ..]
CRUCIAL: Return ONLY the JSON, no other text or explanations!
A.3 Example of Full Evaluation Flow
To make the process concrete, this section details the full evaluation flow for a single sample from the HLE benchmark.
[…]
Step 2: Raw JSON Output from LLM The LLM processes the input and returns only the following raw JSON object, identifying the relevant capability and propensity codes.
  "capab": ["C6"],
  "prop": []
Step 3: Final Enriched Data Record Our pipeline parses the raw JSON and enriches it with metadata to produce the final, structured data record used in our analysis.
  "model": "google/gemini-2.5-flash",
  "id": "hle_194750",
  "benchmark": "hle",
  "evaluation": {
    "capabilities": {
      "Long-horizon planning, forecasting, or strategising": 1,
    "propensities": {}
Appendix B Expert Annotation and Risk Mapping Protocols
This appendix provides a detailed account of the protocols followed by our expert team for two critical stages of this research: (1) the annotation of the gold standard dataset used to validate our LLM evaluator, and (2) the mapping of low-level capabilities and propensities to the high-level systemic risks defined in the EU AI Act Code of Practice.
B.1 Gold Standard Annotation Protocol
The creation of a high-quality, human-annotated gold standard dataset was a prerequisite for validating the LLM-as-judge approach described in Section 3. The following protocol was established to ensure consistency, reliability, and accuracy in this process.
Expert Team Composition
The annotation team consisted of five researchers with complementary expertise spanning technical AI safety, machine learning evaluation methodologies, AI ethics, and regulatory compliance, including specific knowledge of the EU AI Act. This diversity was crucial for interpreting both the technical nuances of the benchmark questions and the legal-ethical intent of the CoP taxonomy.
Annotation Instructions
Each expert was provided with a shared spreadsheet containing the 597 questions of the stratified random sample and the following set of instructions for each question:
  1. 1.
    Thoroughly Review the Sample: Read the complete question, including any provided context, the list of multiple-choice options, and the correct answer. Understanding the full scope of the item is essential for accurate classification.
  2. 2.
    Identify All Relevant Capabilities: Based on the definitions provided in the CoP (and summarized in the main paper), identify all capabilities (C1-C13) that are being evaluated by the question. A capability should be included if it meets any of the classification criteria (Direct Test, Required for Answer, Primary Risk) outlined in Appendix A.
  3. 3.
    Identify All Relevant Propensities: Similarly, identify all propensities (P1-P9) that could be revealed or measured through a model’s response to the question.
  4. 4.
    Provide Rationale for Ambiguity: For any classifications that are not immediately obvious or could be subject to interpretation, provide a brief, written rationale in the designated notes column. This was particularly important for distinguishing between direct and indirect evaluations.
Quality Control and Inter-Rater Reliability
To ensure the quality and consistency of the annotations, we implemented a multi-stage validation process:
  • Calibration Session: Before beginning the main annotation task, the team held a two-hour calibration session. During this session, the team collectively annotated 20 sample questions, discussing interpretations of the CoP definitions and resolving initial ambiguities to establish a shared understanding and consistent "case law."
  • Independent Annotation and Overlap: Each question in the gold standard set was annotated by at least one expert. A randomly selected subset of 20% of the questions (approximately 120 questions) was annotated independently by at least two different experts to allow for the calculation of inter-rater reliability metrics, such as Cohen’s Kappa.
  • Consensus Resolution: After the independent annotation phase, the team met to review all disagreements found in the overlapping set. Each disagreement was discussed until a consensus was reached, and the final classification was recorded. The reasoning for these consensus decisions was documented and used to refine the annotation guidelines for the remainder of the dataset.
B.2 Systemic Risk Mapping Rationale
As noted in Section 4.5, the EU AI Act’s Code of Practice identifies four high-level systemic risks but does not provide an explicit mapping from these risks to the low-level capabilities and propensities. To bridge this interpretive gap for our analysis, our expert team conducted a structured, deliberative exercise to define these "risk pathways." The following rationale details the reasoning behind our mapping for each systemic risk category.
Harmful Manipulation
This risk category encompasses large-scale deception, political interference, and social destabilization. We determined that this risk emerges primarily from the combination of a model’s ability to generate convincing but false content (P3: Hallucination), its potential to target or generate content based on stereotypes (P4: Discriminatory bias), and its functional ability to craft persuasive arguments (C3: Manipulate/persuade/deceive). The risk is amplified by the ability to operate across different data types (C11: Multimodality) for creating deepfakes or other synthetic media, and is fundamentally rooted in a potential divergence from human-well being (P1: Misalignment).
Cyber Offence
This risk relates to significant cyberattacks and infrastructure disruption. The pathway to this risk begins with a foundational knowledge of vulnerabilities and exploits (C1: Offensive cyber). This knowledge becomes actively dangerous when combined with the ability to interact with computer systems and APIs (C12: Tool use) and a tendency to disregard rules (P6: Lawlessness). The risk is significantly magnified if the model possesses a degree of independent decision-making (C4: Autonomy) and exhibits agentic behaviors like pursuing instrumental goals (P7: Power-seeking).
CBRN Risks
This category involves the facilitation of chemical, biological, radiological, or nuclear threats. The primary component is, naturally, a model’s knowledge of these sensitive domains (C2: CBRN capabilities). This knowledge poses a risk when coupled with a propensity to use its capabilities for destructive ends (P2: Harmful deployment) and a disregard for legal or ethical boundaries (P6: Lawlessness). Furthermore, the ability to acquire new skills or information beyond its initial training (C5: Adaptive learning) is a critical risk factor, as it could allow a model to deepen its dangerous knowledge autonomously.
Loss of Control
This risk, central to long-term AI safety concerns, involves autonomous systems operating beyond human oversight and pursuing misaligned goals. We mapped this risk to a cluster of advanced, agentic capabilities. The core components are a high degree of independent decision-making (C4: Autonomy), the ability to reflect on its own state (C7: Self-reasoning), and the capability to actively hide its operations (C8: Evade human oversight). The risk escalates dramatically with the ability to self-modify or create copies (C9: Self-replicate/improve) and to contribute to its own advancement (C10: AI R&D). This entire risk pathway is predicated on a fundamental divergence from human goals (P1: Misalignment) and a potential drive for self-preservation or influence (P7: Power-seeking).